# Rethinking Early-Fusion Strategies for Improved Multispectral Object Detection

### v1

@@ Abstract
Multispectral object detection; feature fusion; weakly supervised learning; knowledge distillationI Introduction
@@
To resolve this conflict, in this work we focus on improving the performance of the early-fusion strategy while maintaining its high efficiency. We first conduct pilot studies and observe that a plain early-fusion strategy cannot consistently obtain improved performances compared to single-modality inputs. Based on this observation, we rethink the early-fusion strategy and summarize three key obstacles: 1) the information interference problem when simply concatenating the multispectral images, 2) the domain gap existing in thermal and RGB images, and 3) the weak feature representation of the single-branch structure. Focusing on these obstacles, we propose corresponding solutions.
@@
Different from previous works, we summarize three key obstacles limiting the early-fusion strategy, including information interference, domain gap, and representation learning.
For each obstacle, we propose the corresponding solution: we develop 1) a ShaPE module to address the information interference issue, 2) a weakly supervised learning method to reduce domain gap and improve semantic localization abilities, and 3) a CoreKD to enhance the representation learning of single-branch networks.
Extensive experiments validate that the early-fusion strategy, equipped with our ShaPE module, weakly supervised learning, and CoreKD technique, shows significant improvement. Additionally, we only retain the ShaPE module during the inference phase. Consequently, our method is efficient and achieves improved performance.
@@ II-A Multispectral Object Detection
According to fusion strategies, multispectral object detection can be classified into three categories: early-fusion, medium-fusion and late-fusion strategies. Previous works [11], [12] and [22] confirm that both medium-fusion and late-fusion strategies outperform the early-fusion strategy.
@@ III Method
Fig. 2 illustrates the overview of our method. We adopt a single-branch structure as the baseline model considering its low memory cost. To boost its performance, we develop three key modules: shape-priority early-fusion (ShaPE), weakly supervised auxiliary learning, and core knowledge distillation (CoreKD). In the following, we describe the ShaPE module in Section III-A, the weakly supervised auxiliary learning method in Section III-B, and the CoreKD in Section III-C. These contributions are designed to enhance information fusion, feature extraction, and feature classification abilities, respectively.
@@ III-B Weakly Supervised Learning Method
To handle this issue, we turn to the powerful Contrastive Language-Image Pre-training (CLIP) [17] model. It has been confirmed that CLIP can bridge domain gaps [18], since it is trained using a huge number of (image, text) pairs. In this context, we feed both RGB and thermal images into the backbone network, and let it learn the representation generated by the CLIP model. Specifically, we first present a CLIP-driven image-level weakly supervised learning method. This method enables the network to recognize the classes of objects in a pair of RGB-T images while locating their coarse regions. For fine-grained localization, we then introduce a box-level weakly supervised learning method. Fig. 6 illustrates the architecture of weakly supervised learning method.
@@
Nevertheless, original CLIP model is only trained for recognizing a single object per image [17] and is not suitable for multi-label classification [36]. To address this issue, we introduce a Divide-and-Aggregation CLIP (DA-CLIP) model. DA-CLIP first divides input images into multiple crops. Each crop is then fed into CLIP. All predictions of these crops are finally aggregated by a max-pooling operation on each class. Considering DA-CLIP may generate inaccurate predictions, we construct a learnable adapter, which consists of three fully-connected (FC) layers, to fine-tune the result of DA-CLIP. To prevent overfitting, we add a dropout layer in the adapter. We denote the predicted probability from the adapter as , where denotes the number of classes.
@@
We visualize attention maps of the backbone network for different classes, as shown in Fig. 8 (b). Using the box-level weak supervision, the backbone network can precisely localize the interest of objects, such as ’Car’. Nevertheless, it may miss some useful information in the image. Therefore, we combine the CLIP-driven image-level weak supervision and the box-level weak supervision. The results presented Fig. 8 (c) show that our weakly supervised learning method can effectively allow the backbone network to localize the important semantic regions.
@@ III-D Loss Function
| Detector | FLOPs (↓) | Parameters (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Car (↑) | Bus (↑) | Motor (↑) | TrafficLight (↑) | Truck (↑) | |
| RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 30.90 | 51.10 | 44.60 | 74.50 | 58.10 | 44.10 | 36.00 | 49.50 |
| Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 29.30 | 47.10 | 59.50 | 71.20 | 55.00 | 35.80 | 10.30 | 50.60 |
| RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 33.50 | 53.70 | 60.20 | 77.30 | 61.70 | 45.50 | 25.80 | 51.80 |
| Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 32.10 | 50.90 | 58.80 | 75.50 | 59.50 | 40.00 | 22.00 | 49.30 |
| + ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 32.90 | 52.30 | 61.00 | 77.20 | 59.30 | 40.40 | 24.40 | 51.50 |
| + ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.50 | 52.70 | 59.70 | 76.60 | 63.10 | 38.60 | 23.70 | 54.70 |
| + ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.70 | 53.10 | 60.90 | 76.50 | 59.30 | 42.90 | 25.80 | 53.30 |
| RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 32.40 | 53.12 | 48.80 | 77.80 | 60.00 | 43.70 | 38.40 | 50.00 |
| Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 29.80 | 48.70 | 64.60 | 73.90 | 54.20 | 36.50 | 15.20 | 47.70 |
| RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 34.60 | 55.30 | 65.20 | 79.80 | 62.60 | 39.50 | 35.20 | 49.80 |
| Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 33.90 | 53.07 | 64.00 | 78.40 | 54.50 | 39.50 | 29.70 | 52.30 |
| + ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.40 | 55.70 | 65.80 | 79.10 | 63.00 | 41.90 | 30.20 | 54.20 |
| + ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.90 | 56.30 | 65.10 | 79.80 | 66.00 | 41.80 | 30.90 | 54.00 |
| + ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 37.10 | 57.60 | 68.50 | 81.30 | 63.30 | 42.70 | 35.90 | 53.90 |
| Detector | FLOPs (↓) | Parameters (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Bicycle (↑) | Car (↑) | |
| RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 28.10 | 59.50 | 45.10 | 55.70 | 77.90 |
| Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 38.50 | 71.00 | 62.40 | 66.30 | 84.30 |
| RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 38.60 | 71.50 | 61.90 | 67.50 | 85.10 |
| Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 37.40 | 69.50 | 60.40 | 63.90 | 84.30 |
| + ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.60 | 71.40 | 61.30 | 68.30 | 84.50 |
| + ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.40 | 71.60 | 61.50 | 68.40 | 85.00 |
| + ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.60 | 71.80 | 61.90 | 68.30 | 85.00 |
| RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 31.90 | 63.70 | 51.70 | 57.60 | 81.80 |
| Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 42.60 | 75.10 | 69.60 | 68.80 | 86.90 |
| RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 42.70 | 75.80 | 69.80 | 70.10 | 87.70 |
| Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 42.20 | 74.70 | 69.70 | 67.50 | 87.00 |
| + ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 42.60 | 75.60 | 69.70 | 70.10 | 87.10 |
| + ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 43.10 | 76.60 | 71.10 | 70.70 | 87.90 |
| + ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 44.00 | 78.10 | 73.10 | 72.40 | 88.80 |

### v2

@@ Abstract
Multispectral object detection; feature fusion; weakly supervised learning; knowledge distillation
I Introduction
@@
The main motivation for this work is to resolve the conflict between detection performance and inference efficiency. To this end, we focus on improving the performance of the early-fusion strategy while maintaining its high computational efficiency. We first conduct pilot studies and observe that a plain early-fusion strategy cannot consistently obtain improved performances compared to single-modality inputs. Based on this observation, we rethink the early-fusion strategy and summarize three key obstacles: 1) the information interference problem when simply concatenating the multispectral images, 2) the domain gap existing in thermal and RGB images, and 3) the weak feature representation of the single-branch structure. Focusing on these obstacles, we propose corresponding solutions.
@@
We systematically analyze the causes of the performance gap between single-branch and two-branch structures. Unlike previous works, we identify and summarize three key obstacles limiting the single-branch early-fusion strategy: information interference, domain gap, and weak feature representation. Notably, information interference between multispectral images is revealed for the first time in this work.
For each obstacle, we propose the corresponding solution: we develop 1) a ShaPE module to address the information interference issue, 2) a weakly supervised learning method to reduce domain gap and improve semantic localization abilities, and 3) a CoreKD to enhance the feature representation of single-branch networks.
Extensive experiments validate that the early-fusion strategy, equipped with our ShaPE module, weakly supervised learning, and CoreKD technique, demonstrates significant improvement. These three modules benefit various common detectors, such as YOLOv5 [2], RetinaNet [22], and GFL [23]. Importantly, only the ShaPE module is retained during the inference phase. Consequently, our method achieves both high performance and efficiency.
@@ II-A Multispectral Object Detection
Multi-source information fusion [24, 25, 26, 27] has exhibited promising application potential in computer vision tasks. In this work, we focus on the multispectral object detection task that uses RGB and thermal image pairs to detect objects. According to fusion strategies, multispectral object detection can be classified into three categories: early-fusion, medium-fusion and late-fusion strategies. Previous works [11], [12] and [28] confirm that both medium-fusion and late-fusion strategies outperform the early-fusion strategy.
@@ III Method
Fig. 2 illustrates the overview of our method, where the training process and the inference process are marked in green and blue, respectively. We adopt a single-branch structure as the baseline model considering its low memory cost. To boost its performance, we develop three key modules: shape-priority early-fusion (ShaPE), weakly supervised auxiliary learning, and core knowledge distillation (CoreKD). Note that only the training process requires weakly supervised auxiliary learning and CoreKD, and both are removed during the inference phase. Consequently, our method adds only the ShaPE module to the early-fusion single-branch structure during the inference phase. In the following sections, we describe the ShaPE module in Section III-A, the weakly supervised auxiliary learning method in Section III-B, and CoreKD in Section III-C.
@@ III-B Weakly Supervised Learning Method
To handle this issue, we turn to the powerful Contrastive Language-Image Pre-training (CLIP) [17] model. It has been confirmed that CLIP can bridge domain gaps [18, 40, 41, 42, 43], since it is trained using a huge number of (image, text) pairs. In this context, we feed both RGB and thermal images into the backbone network, and let it learn the representation generated by the CLIP model. Specifically, we first present a CLIP-driven image-level weakly supervised learning method. This method enables the network to recognize the classes of objects in a pair of RGB-T images while locating their coarse regions. For fine-grained localization, we then introduce a box-level weakly supervised learning method. Fig. 6 illustrates the architecture of weakly supervised learning method.
@@
Nevertheless, original CLIP model is only trained for recognizing a single object per image [17] and is not suitable for multi-label classification [44, 45]. To address this issue, we introduce a Divide-and-Aggregation CLIP (DA-CLIP) model. DA-CLIP first divides input images into multiple crops. Each crop is then fed into CLIP. All predictions of these crops are finally aggregated by a max-pooling operation on each class. Considering DA-CLIP may generate inaccurate predictions, we construct a learnable adapter, which consists of three fully-connected (FC) layers, to fine-tune the result of DA-CLIP. To prevent overfitting, we add a dropout layer in the adapter. We denote the predicted probability from the adapter as , where denotes the number of classes.
@@
We visualize attention maps of the backbone network for different classes, as shown in Fig. 8 (b). Using the box-level weak supervision, the backbone network can precisely localize the interest of objects, such as ’Car’. Nevertheless, it may miss some useful information in the image. Therefore, we combine the CLIP-driven image-level weak supervision and the box-level weak supervision. The results presented in Fig. 8 (c) show that our weakly supervised learning method can effectively allow the backbone network to localize the important semantic regions.
@@ III-C Core Knowledge Distillation
Mathematical Foundation of CoreKD. We first explain the mathematical foundation of traditional feature distillation and analyze its weaknesses. Then, we introduce the mathematical foundation of our CoreKD. Finally, we compare the results of our CoreKD with those of the traditional one.
Traditional feature distillation uses a convolution layer to align feature channel numbers between student and teacher models. We denote the convolution layer as a function in the Eq. (6). Next, we denote the weighting factor of as . This function is used to convert an input feature into an output feature , i.e., . We denote the vector at an arbitrary spatial location of as , the ith row vector in the weighting factor as , and the corresponding value in as . The mathematical relation between , , and can be represented as
where the operator ‘’ indicates a dot product. The dot product computation can be viewed as the projection of the vector onto the vector , as shown in Fig. 11 (a). We can infer that the generation of is related to the and but has no relation to the teacher’s features. This implies that the traditional feature distillation merely focuses on enforcing the student to learn all information from the teacher without considering whether the teacher’s features are beneficial to downstream tasks.
On the contrary, our CoreKD uses the weighting factor of the teacher model to align feature channel numbers between the student and teacher models, as shown in Eq. (8). We denote the ith row vector of as and the vector at an arbitrary spatial location of as . Then we can write the ith loss value of as
This loss value calculation process is illustrated in Fig. 11 (b). From the above analyses, we have two key observations: 1) our CoreKD projects the student and teacher features into an identical space constructed by , and 2) our CoreKD does not enforce the student feature to be the same as the teacher feature but rather focuses on minimizing the projected distances. Since the values within the weighting factor reflect the importance of teacher features, our CoreKD enables the student model to learn beneficial information for downstream tasks from the teacher model.
Since the experimental results involve the introduction of both datasets and implementation details, we arrange the comparison results of our CoreKD with the traditional feature distillation in the experimental section. For details, please refer to Section IV-C.
@@ III-D Loss Function
IV Experiments
IV-A Experimental Setup
Datasets. Our experiments are conducted on the M3FD dataset [1] and FLIR dataset [39]. M3FD dataset contains 4200 pairs of RGB and thermal images. These image pairs are well aligned. The dataset contains 6 classes of objects: ‘Person’, ‘Car’, ‘Bus’, ‘Motorcycle’, ‘Traffic Light’, and ‘Truck’. Since this dataset doesn’t provide unified data splits, previous works have used a random splitting approach to determine the train and validation sets [1]. However, images in this dataset are sampled from video sequences, meaning that two adjacent frames may contain identical content. In this context, the random splitting approach results in information leakage between the train and validation sets. To address this problem, we first manually divide the dataset into 73 video segments based on different scenes. Then, we collect the first 70% of images in each video segment as the train set and the remaining images as the validation set. Finally, we obtain 2905 and 1295 pairs of RGB-T images in the train and validation sets, respectively. We name this data split ‘M3FD-zxSplit’ and release it to the public111https://github.com/XueZ-phd/Efficient-RGB-T-Early-Fusion-Detection. For the performance evaluation in Section IV-B, we use this data split. When comparing with state-of-the-art approaches in Section IV-D, we employ both ‘M3FD-zxSplit’ and random splitting. Our random splitting refers to randomly selecting 80% images as the train set and the remaining images as the validation set. FLIR dataset originally contains unaligned RGB-T image pairs. The work [49] develops a data-processing approach to align these images and obtain 7381, 1056, and 2111 image pairs in the train, validation, and test sets. This dataset contains 3 classes: ‘Person’, ‘Bicycle’ and ‘Car’.
Inference Efficiency Evaluations. We assess the inference efficiency of our method (Python implementation) on the edge device NVIDIA AGX Orin with 64GB memory. We also evaluate the complexity of our method using FLOPs and the number of parameters. All results are presented in Tables I and II.
Implementation Details. We incorporate our three key modules into commonly-used one-stage detectors, including RetinaNet [22], GFL [23], and YOLOv5 [2]. For RetinaNet and GFL, we adopt the implementations in MMDetection toolbox [50]. For YOLOv5, we use its official implemtation [2].
In the early-fusion strategy based on RetinaNet [22] and GFL [23] detectors, we use ResNet-50 [48] as the backbone network. For a fair comparison, we use the same backbone network in the medium-fusion strategy. Notably, in the CoreKD technique, the teacher model utilizes ResNet-101 [48] as the backbone network. For both strategies, we train for 12 epochs using the SGD optimizer with a batch size of 4. The initial learning rate is set to 0.01 and is decayed by 0.1 at epochs 8 and 11. Random horizontal flipping is employed as a data augmentation technique.
For the early-fusion strategy based on YOLOv5, we use YOLOv5-small as the baseline detector, and we use YOLOv5-large to construct a teacher model in the CoreKD technique. For both strategies, we train for 36 epochs with a batch size of 16. We keep all other hyperparameters consistent with the official settings of the YOLOv5 repository [2].
To standardize data for RetinaNet and GFL detectors, we calculate the mean value and standard deviation of RGB and thermal images for the M3FD and FLIR datasets. All experiments use the 640 512 image resolution. For the M3FD dataset, we obtain meanrgb = [128.2, 129.3, 125.3], stdrgb = [49.1, 50.2, 53.5], meant = [84.1, 84.1, 84.1], and stdt = [50.6, 50.6, 50.6]. For the FLIR dataset, we obtain meanrgb = [149.4, 148.7, 141.7], stdrgb = [49.3, 52.8, 59.0], meant = [135.7, 135.7, 135.7], and stdt = [63.6, 63.6, 63.6]. For the YOLOv5 detector, we normalize both RGB and thermal images to the range of [0, 1] following its official implementations.
@@ IV-A Experimental Setup
Detector | FLOPs (↓) | Parameter (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Car (↑) | Bus (↑) | Motor (↑) | TrafficLight (↑) | Truck (↑) | |
RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 31.03±0.09 | 51.30±0.16 | 44.57±0.12 | 74.87±0.26 | 57.80±0.24 | 44.30±0.22 | 36.63±0.58 | 49.70±0.16 |
Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 29.27±0.05 | 46.97±0.09 | 59.17±0.34 | 71.17±0.29 | 54.17±0.69 | 35.90±0.22 | 10.43±0.09 | 50.83±0.21 |
RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 33.43±0.05 | 53.63±0.05 | 60.10±0.08 | 77.27±0.05 | 61.63±0.05 | 45.43±0.09 | 25.67±0.12 | 51.80±0.00 |
Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 32.03±0.05 | 50.70±0.16 | 58.93±0.26 | 75.37±0.09 | 58.97±0.39 | 39.97±0.21 | 21.67±0.47 | 49.33±0.05 |
+ ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 32.80±0.08 | 51.97±0.24 | 60.20±0.59 | 77.10±0.08 | 58.83±0.33 | 39.77±0.58 | 24.57±0.54 | 51.33±0.17 |
+ ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.53±0.05 | 52.90±0.16 | 59.10±0.71 | 77.07±0.41 | 62.00±0.78 | 40.97±1.67 | 24.03±0.85 | 54.30±0.71 |
+ ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 33.53±0.17 | 53.23±0.09 | 61.47±0.49 | 76.40±0.08 | 59.20±0.08 | 43.10±0.22 | 25.97±0.12 | 53.17±0.19 |
RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 32.23±0.12 | 53.10±0.16 | 48.67±0.12 | 77.43±0.26 | 60.27±0.60 | 43.50±0.16 | 39.07±0.94 | 49.63±0.39 |
Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 29.50±0.22 | 48.27±0.42 | 64.27±0.34 | 73.73±0.12 | 52.50±1.99 | 36.50±0.0 | 15.10±0.14 | 47.37±0.25 |
RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 34.17±0.33 | 54.47±0.60 | 65.37±0.12 | 79.83±0.05 | 61.20±1.00 | 37.00±1.79 | 34.80±0.57 | 48.73±0.77 |
Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 33.50±0.28 | 52.77±0.25 | 64.03±0.12 | 78.20±0.16 | 53.00±1.31 | 39.27±0.26 | 30.47±0.56 | 51.57±0.54 |
+ ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.17±0.17 | 55.50±0.22 | 65.80±0.24 | 79.10±0.08 | 62.33±1.24 | 41.33±0.42 | 30.80±0.42 | 53.67±0.41 |
+ ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 35.23±0.53 | 55.97±0.24 | 65.73±0.49 | 79.57±0.17 | 60.50±4.02 | 43.10±1.06 | 33.73±2.01 | 53.10±1.94 |
+ ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 37.03±0.09 | 57.70±0.08 | 68.43±0.25 | 81.23±0.09 | 63.37±0.33 | 43.90±0.86 | 35.77±0.12 | 53.53±0.29 |
Detector | FLOPs (↓) | Parameter (↓) | Time (↓) | mAP (↑) | mAP50 (↑) | Person (↑) | Bicycle (↑) | Car (↑) | |
RGB | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 28.10±0.0 | 59.47±0.12 | 44.93±0.39 | 55.70±0.08 | 77.80±0.08 |
Thermal | RetinaNet-Res50 | 61.893G | 36.434M | 0.106s | 35.53±0.05 | 70.93±0.05 | 62.17±0.17 | 66.37±0.09 | 84.27±0.05 |
RGB-T Medium Fusion | RetinaNet-Res50 | 94.611G | 47.582M | 0.170s | 38.50±0.08 | 71.57±0.05 | 61.93±0.29 | 67.60±0.14 | 85.17±0.09 |
Baseline: RGB-T Early Fusion | RetinaNet-Res50 | 62.164G | 36.434M | 0.110s | 37.47±0.05 | 69.57±0.05 | 60.70±0.22 | 63.77±0.09 | 84.37±0.05 |
+ ShaPE | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.70±0.14 | 71.60±0.22 | 61.40±0.54 | 68.60±0.36 | 84.70±0.22 |
+ ShaPE + WeakSup. | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.80±0.29 | 72.07±0.37 | 62.50±0.96 | 68.77±0.97 | 85.03±0.21 |
+ ShaPE + WeakSup. + CoreKD | RetinaNet-Res50 | 62.218G | 36.434M | 0.149s | 38.83±0.17 | 72.23±0.31 | 62.27±0.26 | 69.23±0.66 | 85.10±0.22 |
RGB | GFL-Res50 | 61.392G | 32.270M | 0.110s | 31.73±0.12 | 63.77±0.05 | 51.70±0.08 | 57.87±0.25 | 81.77±0.05 |
Thermal | GFL-Res50 | 61.392G | 32.270M | 0.110s | 42.40±0.14 | 75.07±0.05 | 69.80±0.16 | 68.47±0.24 | 86.93±0.05 |
RGB-T Medium Fusion | GFL-Res50 | 94.110G | 43.419M | 0.172s | 42.60±0.08 | 76.07±0.21 | 70.07±0.19 | 70.57±0.41 | 87.63±0.05 |
Baseline: RGB-T Early Fusion | GFL-Res50 | 61.663G | 32.271M | 0.114s | 41.90±0.22 | 74.77±0.17 | 69.70±0.33 | 67.73±0.26 | 87.00±0.00 |
+ ShaPE | GFL-Res50 | 61.718G | 32.271M | 0.151s | 42.40±0.16 | 75.77±0.17 | 69.97±0.21 | 70.13±0.37 | 87.23±0.09 |
+ ShaPE + WeakSup. | GFL-Res50 | 61.718G | 32.271M | 0.151s | 42.93±0.24 | 76.90±0.22 | 71.30±0.14 | 71.40±0.50 | 87.97±0.09 |
+ ShaPE + WeakSup. + CoreKD | GFL-Res50 | 61.718G | 32.271M | 0.151s | 44.00±0.00 | 78.17±0.05 | 73.03±0.05 | 72.63±0.17 | 88.80±0.00 |
IV-B Performance Evaluation of Proposed Modules
Table I and Table II present the performance of our method on the M3FD [1] and FLIR [39] datasets. Key observations include: (1) the medium-fusion strategy adds more parameters and FLOPs compared to the early-fusion strategy; (2) the medium-fusion strategy achieves better performance compared to single-modality inputs, whereas the plain early-fusion strategy does not consistently improve performance; (3) our EME method, incorporating the ShaPE module, weakly supervised learning, and CoreKD techniques into the plain early-fusion strategy, achieves significant performance improvement without significantly increasing parameters and FLOPs; (4) the inference time of our EME method is longer than that of the baseline method, since the structural similarity computation process has not been optimized when calculating the self-gating mask; (5) our EME method can outperform the medium-fusion strategy in both performance and efficiency to some extent; (6) Both architectures: “Baseline + ShaPE + WeakSup.” and “Baseline + ShaPE + WeakSup. + CoreKD” have the same FLOPs, parameters, and inference time as “Baseline + ShaPE”. This is because both the weakly supervised learning method and CoreKD are removed in the inference phase, while only the ShaPE module is retained.
IV-C Performance Evaluation of Variants
Feature Distillation. We compare the results of traditional feature distillation with those of our CoreKD. Table III and Table IV present comparison results on the M3FD and FLIR datasets, respectively. For comprehensive comparisons, we adopt RetinaNet-Res50 and GFL-Res50 as baseline detectors in these two tables. From the results, we can observe that our CoreKD consistently achieves superior performance compared to traditional feature distillation. For example, our CoreKD (72.23%) obtains a 1.26% mAP50 absolute gain over the traditional one (70.97%) when using RetinaNet-Res50 on the FLIR dataset.
Method | Detector | mAP (↑) | mAP50 (↑) | ||
RetinaNet-Res50 | 32.03±0.05 | 50.70±0.16 | ||
RetinaNet-Res50 | 32.17±0.12 | 52.43±0.05 | ||
RetinaNet-Res50 | 33.53±0.17 | 53.23±0.09 | ||
GFL-Res50 | 33.50±0.28 | 52.77±0.25 | ||
GFL-Res50 | 35.83±0.17 | 57.07±0.09 | ||
GFL-Res50 | 37.03±0.09 | 57.70±0.08 |
Method | Detector | mAP (↑) | mAP50 (↑) | ||
RetinaNet-Res50 | 37.47±0.05 | 69.57±0.05 | ||
RetinaNet-Res50 | 38.10±0.14 | 70.97±0.12 | ||
RetinaNet-Res50 | 38.83±0.17 | 72.23±0.31 | ||
GFL-Res50 | 41.90±0.22 | 74.77±0.17 | ||
GFL-Res50 | 43.63±0.09 | 77.80±0.08 | ||
GFL-Res50 | 44.00±0.00 | 78.17±0.05 |
Backbone Network. We evaluate our EME method using ResNet-101 [48] as the backbone network on the M3FD and FLIR datasets, and present the results in Table V. We observe that detectors using ResNet-101 consistently achieve better performance than those using ResNet-50. For example, RetinaNet with ResNet-101 (74.87%) obtains a 2.64% mAP50 absolute gain over ResNet-50 (72.23%).
Datasets | Detector | Backbone | FLOPs | mAP (↑) | mAP50 (↑) |
M3FD | RetinaNet | ResNet-50 | 62.218G | 33.53±0.17 | 53.23±0.09 |
RetinaNet | ResNet-101 | 65.392G | 34.23±0.12 | 54.63±0.45 | |
GFL | ResNet-50 | 61.718G | 37.03±0.09 | 57.70±0.08 | |
GFL | ResNet-101 | 64.892G | 37.37±0.12 | 58.60±0.08 | |
FLIR | RetinaNet | ResNet-50 | 62.218G | 38.83±0.17 | 72.23±0.31 |
RetinaNet | ResNet-101 | 65.392G | 40.67±0.05 | 74.87±0.34 | |
GFL | ResNet-50 | 61.718G | 44.00±0.00 | 78.17±0.05 | |
GFL | ResNet-101 | 64.892G | 44.47±0.05 | 79.57±0.05 |